OpenStudy (josee):

Find all values of x at which the tangent line to the graph of y = x^2/(x + 3) is horizontal.

OpenStudy (anonymous):

For Horizontal \[\frac{dy}{dx} = 0 \] \[\frac{2x(x+3)- x^2*1}{(x+3)^2} = 0\] \[2x^2 - x^2 + 6x = 0\] \[x^2 + 6x= 0\] \[x(x + 6) = 0 \] \[x = 0,-6\] At x = 0 , -6 y must be 0 and \(\frac{36}{-3}\) respectively . Hence our ordered pairs are (0,0) and \((-6, -12)\)